# Gravitational signatures of a nonlinear electrodynamics in $f(R,T)$ gravity

### v1

@@ VIII Time–domain solution: fermionic case
Up to now, we have examined the time evolution of scalar, vector, and tensor perturbations. To complete this analysis, we now turn our attention to the time–domain behavior of spinorial perturbations. In this manner, Fig. 22 shows the time evolution of the spinor perturbation $\tilde{\psi}$ for fixed $\alpha=\beta=-0.01$ and varying $Q=0.6$, $0.7$, $0.8$, and $0.9$. The left and right panels display results for $\ell=1$ and $\ell=2$, respectively, highlighting the gradual damping of the signal. In Fig. 23, the decay is further analyzed through a $\ln$ plot of $\ln|\tilde{\psi}|$ versus $t$. Finally, Fig. 24 presents a $\ln–\ln$ plot of $\ln|\tilde{\psi}|$ against $\ln|t|$, confirming the emergence of power–law tails.
@@
Figure 23: Logarithmic representation of spinorial perturbations, displaying $\ln|\tilde{\psi}|$ versus time $t$ for fixed $\alpha=\beta=-0.01$ and varying $Q=0.6$, $0.7$, $0.8$, and $0.9$. The left and right panels illustrate the behavior for angular momentum modes $\ell=1$ and $\ell=2$, respectively.
Figure 24: Late–time dynamics of spinor perturbations illustrated using a $\ln$$\ln$ plot of $\ln|\tilde{\psi}|$ against $\ln|t|$, with deformation parameters set to $\alpha=\beta=-0.01$ and charge values $Q=0.6$, $0.7$, $0.8$, and $0.9$. Results for $\ell=1$ and $\ell=2$ are shown in the left and right panels, respectively.
@@ X Strong field lensing regime
From this point onward, all quantities labeled with the subscript “00” refer to their values at the radial coordinate $r=r_{0}$. When examining the propagation of a single photon, it is convenient—and physically justified—to restrict the impact parameter $b$ to positive values. Since $b$ is conserved along the photon’s path, it admits the following representation:
@@ XI.1 Topological thermodynamics
Recent advances in black hole thermodynamics have introduced a topological perspective as a powerful framework for analyzing critical points and phase transitions wei2022black ; wu2025novel ; yerra2022topology ; wu2023topological ; wu2023topological1 ; fang2023revisiting ; zhang2023bulk ; gogoi2023thermodynamic ; fan2023topological ; sadeghi2024thermodynamic . The main idea comes from Duan’s $\phi$–mapping topological current theory Duane1984 , which systematically assigns topological charges to critical points–where the phase transitions occur.
@@
$n^{r_{h}}=\frac{\phi^{r_{h}}}{||\phi||},\quad n^{\theta}=\frac{\phi^{\theta}}{||\phi||}.$ (126)
The normalized vector field is illustrated in Fig. 28, where the critical point is precisely identified as the zero of the vector field. In this case, $r_{C_{1}}=0.472473$ and $r_{C_{2}}=0.834131$ mark the location of the critical points. Two closed contours enclose these points to examine their topological characteristics, through which the topological charges of $+1$ and $-1$ correspond to critical points $C_{1}$ and $C_{2}$, respectively. Based on the classification introduced in Ref. wei2022black , this result indicates that the critical points are of the novel and conventional type, respectively.
@@ XI.2 Topological photon sphere
$H(r,\theta)$ $=\sqrt{\frac{-\mathrm{A}(r)}{\mathrm{D}(r)}}$ (127)
@@
Photon spheres can be interpreted as topological defects that emerge at points where the vector field vanishes. Following Duane’s topological approach Duane1984 , when a closed contour encircles such a zero point, the resulting net topological charge corresponds to the winding number around that loop. The winding number is obtained by considering a closed, smooth, positively oriented loop $C_{i}$ enclosing the $i$–th critical point. Then, the winding number is calculated by
@@
Thus, each photon sphere can be characterized by a distinct topological charge, which takes discrete values of $\mathcal{Q}=\pm 1$. According to the classification introduced in Wei2020 ; Sadeghi2024 , a photon sphere with topological charge $\mathcal{Q}=-1$ is associated with unstable, while one with topological charge $\mathcal{Q}=+1$ is associated with stability. It is worth mentioning that when a closed contour does not include any zero points of the vector field, the associated topological charge is zero, signifying that no photon spheres are present within the region. The vector space of photon sphere field is illustrated in Fig. 30 for fixed values of $M=1$, $Q=0.5$ and $\alpha=\beta=-0.1$. Figure 30 displays a critical point situated outside the event horizon, indicated by a dot at the photonic radius $r_{ph}=2.82286$. This point possesses a topological charge of $-1$, which characterizes it as an unstable photonic radius.
@@ XII Conclusion
This study focused on exploring the gravitational signatures associated with a specific solution of modified electrodynamics in the framework of $f(R,T)$ gravity, as proposed in Ref. Rois:2024iiu . We began by commenting on the derivation of the black hole solution and examined the conditions under which the event horizon $r_{h}$ remains both real and positive. The analysis revealed two important constraints: $02\alpha\beta Q^{4}-Q^{4}>$ 0 and $M>Q$.
@@ Acknowledgments
A. A. Araújo Filho is supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) and Fundação de Apoio à Pesquisa do Estado da Paraíba (FAPESQ), project No. 150891/2023-7. Also, V. B. Bezerra is partially supported by the Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) grant number 307211/2020-7.

### v2

@@ VIII Time–domain solution: fermionic case
[…]
@@
[…]
@@ X Strong field lensing regime
From this point onward, all quantities labeled with the subscript “0” refer to their values at the radial coordinate $r=r_{0}$. When examining the propagation of a single photon, it is convenient—and physically justified—to restrict the impact parameter $b$ to positive values. Since $b$ is conserved along the photon’s path, it admits the following representation:
@@ XI.1 Topological thermodynamics
Recent advances in black hole thermodynamics have introduced a topological perspective as a powerful framework for analyzing critical points and phase transitions wei2022black ; wu2025novel ; yerra2022topology ; wu2023topological ; wu2023topological1 ; fang2023revisiting ; zhang2023bulk ; gogoi2023thermodynamic ; fan2023topological ; sadeghi2024thermodynamic ; Wu:2023sue ; Wu:2023fcw ; Zhu:2024zcl ; Afshar:2024bgi . The main idea comes from Duan’s $\phi$–mapping topological current theory Duane1984 , which systematically assigns topological charges to critical points–where the phase transitions occur.
@@
[…]
The normalized vector field is illustrated in Fig. 29, where the critical point is precisely identified as the zero of the vector field. In this case, $r_{C_{1}}=0.472473$ and $r_{C_{2}}=0.834131$ mark the location of the critical points. Two closed contours enclose these points to examine their topological characteristics, through which the topological charges of +1+1 and 1-1 correspond to critical points $C_{1}$ and $C_{2}$, respectively. Based on the classification introduced in Ref. wei2022black , this result indicates that the critical points are of the novel and conventional type, respectively.
@@ XI.2 Topological photon sphere
$H(r,\theta)$ $=\sqrt{\frac{\mathrm{A}(r)}{\mathrm{D}(r)}}$ (127)
@@
Photon spheres can be interpreted as topological defects that emerge at points where the vector field vanishes. Following Duane’s topological approach Duane1984 , when a closed contour encircles such a zero point, the resulting net topological charge corresponds to the winding number around that loop. The winding number is obtained by considering a closed, smooth, positively oriented loop $C_{i}$ enclosing the ii–th critical point. Then, the winding number is calculated by
@@
Thus, each photon sphere can be characterized by a distinct topological charge, which takes discrete values of $\mathcal{Q}=\pm 1$. According to the classification introduced in Wei2020 ; Sadeghi2024 , a photon sphere with topological charge $\mathcal{Q}=-1$ is associated with unstable, while one with topological charge $\mathcal{Q}=+1$ is associated with stability. It is worth mentioning that when a closed contour does not include any zero points of the vector field, the associated topological charge is zero, signifying that no photon spheres are present within the region. The vector space of photon sphere field is illustrated in Fig. 31 for fixed values of $M=1$, $Q=0.5$ and $\alpha=\beta=-0.1$. Figure 31 displays a critical point situated outside the event horizon, indicated by a dot at the photonic radius $r_{ph}=2.82286$. This point possesses a topological charge of 1-1, which characterizes it as an unstable photonic radius.
@@ XII Conclusion
This study focused on exploring the gravitational signatures associated with a specific solution of modified electrodynamics in the framework of $f(R,T)$ gravity, as proposed in Ref. Rois:2024iiu . We began by commenting on the derivation of the black hole solution and examined the conditions under which the event horizon $r_{h}$ remains both real and positive. The analysis revealed two important constraints: $2\alpha\beta Q^{4}-Q^{4}>0$ and $M>Q$.
@@ Acknowledgments
A. A. Araújo Filho is supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) and Fundação de Apoio à Pesquisa do Estado da Paraíba (FAPESQ), project No. 150891/2023-7. Also, V. B. Bezerra is partially supported by the Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) grant number 307211/2020-7. I. P. L. was partially supported by the National Council for Scientific and Technological Development - CNPq, grant 312547/2023-4. I. P. L. would like to acknowledge networking support by the COST Action BridgeQG (CA23130) and the COST Action RQI (CA23115), supported by COST (European Cooperation in Science and Technology). N. H. would like to acknowledge the contribution of the COST Action CA21106 - COSMIC WISPers in the Dark Universe: Theory, astrophysics and experiments (CosmicWISPers), the COST Action CA21136 - Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse), the COST Action CA23130 - Bridging high and low energies in search of quantum gravity (BridgeQG).